# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

### v1

@@ Abstract
Machine Learning (ML) in low-data settings remains an underappreciated yet crucial problem. This challenge is pronounced in low-to-middle income countries where access to large datasets is often limited or even absent. Hence, data augmentation methods to increase the sample size of datasets needed for ML are key to unlocking the transformative potential of ML in data-deprived regions and domains. Unfortunately, the limited training set constrains traditional tabular synthetic data generators in their ability to generate a large and diverse augmented dataset needed for ML tasks. To address this technical challenge, we introduce CLLM, which leverages the prior knowledge of Large Language Models (LLMs) for data augmentation in the low-data regime. While diverse, not all the data generated by LLMs will help increase utility for a downstream task, as for any generative model. Consequently, we introduce a principled curation process, leveraging learning dynamics, coupled with confidence and uncertainty metrics, to obtain a high-quality dataset. Empirically, on multiple real-world datasets, we demonstrate the superior performance of LLMs in the low-data regime compared to conventional generators. We further show our curation mechanism improves the downstream performance for all generators, including LLMs. Additionally, we provide insights and understanding into the LLM generation and curation mechanism, shedding light on the features that enable them to output high-quality augmented datasets. CLLM  paves the way for wider usage of ML in data scarce domains and regions, by allying the strengths of LLMs with a robust data-centric approach.
@@ 1 Introduction
No data, No Machine Learning. Machine learning (ML) has transformed numerous industries, but its wider adoption is hindered by a pervasive roadblock: insufficient data. Specifically, the use of ML algorithms presumes the availability and access to large datasets for training, be it in the form of labeled or unlabeled data. Unfortunately, real-world domains are often data scarce: (i) in healthcare and finance, collecting annotations can be expensive or practically impossible; (ii) in developing and low-to-middle income countries (LMICs), digital infrastructure (such as electronic healthcare records (EHRs)) can be limited or nonexistent (Ade-Ibijola and Okonkwo, 2023; Asiedu et al., 2023; Owoyemi et al., 2020; Mollura et al., 2020; Alami et al., 2020; Ciecierski-Holmes et al., 2022) and (iii) within large datasets, there can be (ethnic) minorities that are underrepresented. This lack of data has serious consequences: to sideline these settings to the peripheries of ML advancements and prevent the development of accurate models. How can we build a reliable ML model in this low-data regime, where we have so few samples? Solving this problem is a major opportunity that would unlock the potential of ML across society, domains, and regions.
Aim. To address this important yet undervalued low-data problem, we aim to augment the small labeled dataset ($n<100$) with synthetic samples. We focus on tabular data, as defining augmentations is non-trivial and can easily result in nonsensical or invalid samples. Moreover, tabular domains like healthcare (of value in LMICs) are often where data scarcity is acute.
@@
Related work. Data augmentation is a widely used and different approach to address data scarcity in tabular data contexts. Methods are either based on generative models (Ghosheh et al., 2023; Biswas et al., 2023; Wang and Pai, 2023; Machado et al., 2022; Tanaka and Aranha, 2019) such as GANs (Xu et al., 2019), VAEs (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), Score-based models (Kotelnikov et al., 2022; Kim et al., 2022), or alternatively traditional methods such as SMOTE (Chawla et al., 2002; Wang and Pai, 2023; Machado et al., 2022). However, in ultra low-data regimes ($n<100$), the training data may not describe the full data distribution well, despite it being i.i.d. draws. Consequently, this harms conventional methods since the augmented data may not be sufficiently diverse and accurate, restricting the generalizability of predictive models trained on such data. Tangentially, prior works have tackled data scarcity in the tabular setting via the lens of transfer learning, where prior knowledge can be transferred from a pretrained model (Levin et al., 2022; Jin and Ucar, 2023) or a knowledge graph (Margeloiu et al., 2022; Ruiz et al., 2023), which might not be available in all settings. Recent work has shown the potential of fine-tuning Large Language Models (LLMs) for tabular data generation (Borisov et al., 2023). While LLMs offer some degree of prior knowledge, there are two challenges in our setting. First, it is computationally expensive to fine-tune LLMs, while needing specialized hardware —luxuries often not available in LMICs, thereby limiting applicability in such settings. Second, fine-tuning often assumes a large number of samples. In our low-data setting it could lead to overfitting and low-quality generated samples, and hence poor downstream models—as we show for Borisov et al. (2023) in Sec. 3.
@@
Curated LLMs. To address these challenges, we propose Curated LLM (CLLM). First, CLLM  leverages the in-context capabilities of LLMs for generation, thereby reducing the computational burden. We also posit for the low-data regime; the diverse pretraining corpus of LLMs carries valuable prior knowledge, which may offer more diversity in their generation compared to other conventional tabular generators. Of course, LLMs are not perfect. Balancing the utility of LLMs against the risk of noisy, irrelevant data is important for downstream performance, hence requiring systemic assessment of the generated data. In fact, this issue is vital for any generative model.
@@
This motivates the second key aspect of CLLM, i.e. post-generation data curation. This addresses the overlooked aspect that not all of the synthetic samples are useful to downstream model performance, with some samples even harmful. We anchor our approach with ideas from learning theory that show the behavior of individual data samples during training, called learning dynamics, provides a salient signal about the value of samples to a learner (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). To provide intuition, samples with variable predictions might be considered ambiguous or other samples might never be learned correctly and could harm a model. In CLLM, we study the learning dynamics of the synthetic data samples, with respect to a model trained on the small real dataset. We then analyze these dynamics by computing two key metrics: confidence and aleatoric (data) uncertainty. These metrics form the basis for curating the synthetic samples. We aim to enable a highly performant downstream model when trained on the curated dataset.
@@
Contributions: CLLM  is a novel data augmentation approach allying the strengths of LLMs with a robust data curation mechanism to improve data augmentation in the ultra low-data regime $(n<100)$, bringing several contributions: 1⃝ Improved performance: we empirically demonstrate on 7 real-world datasets that CLLM  enables superior downstream performance compared to 6 widely used tabular data generative models and data augmentation techniques. 2⃝ Value of curation: we show the overlooked aspect of synthetic data curation improves downstream performance across the generative models. This highlights the flexibility and broad utility of our curation mechanism for data augmentation. 3⃝ Insights: we dissect the two aspects of CLLM (LLM and data curation) along a variety of dimensions, providing insights and understanding into why the approach is beneficial. We show the largest gains are for underrepresented subgroups and in ultra low-data settings. These contributions pave the way towards wider usage of ML across society, domains and regions.
Ethical considerations. LLMs may make errors and may reflect or exacerbate societal biases that are present in their data (Li et al., 2023). Though the curation in CLLM improves synthetic data quality, it does not directly aim to remove biases. The quality and fairness of generated data should always be evaluated. More research into LLM bias is required before methods like CLLM should be applied to real-world sensitive settings like healthcare and finance.
@@ 2 CLLM: Synergy of LLM Generation and Data Curation
Set-up. Given feature space $\mathcal{X}$, and label space $\mathcal{Y}=\{1,...,k\}$, we assume that we only have a small labeled dataset $D_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, with $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ and $n<100$ (ultra-low data setting). Assume $D_{\mathrm{train}}$ is drawn i.i.d. from the real distribution $p_{R}(X,Y)$. We also assume access to a pretrained LLM to generate samples. We denote the output distribution of the LLM as $p_{\Phi}(X,Y)$, with $\Phi$ containing parameters that we control (e.g., input prompts). Our goal is to generate a dataset to augment the small $D_{\mathrm{train}}$, and subsequently use it to train a classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$. Successful augmentation will provide a better classifier $f$, than if we had trained $f$ on the small $D_{\mathrm{train}}$ itself. We measure downstream performance on a separate held-out dataset of real data, $D_{\mathrm{test}}$.
@@
In each sub-section we describe and motivate the design of the different aspects of CLLM  (LLM and curation). Furthermore, we provide insights and understanding into their role in improving data utility, which we later quantify on multiple real-world datasets in Sec. 3.
@@ 2.1 Data generation with LLMs based on a small $D_{\mathrm{train}}$
As outlined in Sec. 1, in the ultra low-data regime, conventional tabular generative models (e.g. CTGAN, TVAE) are constrained by the limited $D_{\mathrm{train}}$ and may not generate sufficiently diverse and/or accurate synthetic data. To address this challenge, we propose to leverage LLMs, building on their large-scale pretraining. We first outline the desirable features of LLMs for tabular data generation when we have very few samples, then describe design choices to satisfy these.
  • Prior knowledge. LLMs have been pretrained with a vast corpus of information (Chowdhery et al., 2022; Singhal et al., 2023). When prompted to generate samples with limited real data, LLMs can leverage this encoded prior information about similar problems and feature-label relationships to enhance both accuracy and diversity of generation.
@@
Motivation for a frozen LLM. Using a frozen black-box LLM (e.g. GPT-4 or GPT-3.5) is computationally cheaper and requires less specialized hardware (i.e. GPUs) compared to fine-tuning. This relates to settings described in Sec. 1, such as LMICs, where we may not have the computational resources to fine-tune an LLM. Even in settings where fine-tuning is possible, we show empirically in Sec. 3 that LLM fine-tuning (e.g. GReaT baseline) is suboptimal in ultra-low data settings ($n<100$) compared to providing in-context examples coupled with curation.
Dissecting the LLM’s generative features. We now investigate various dimensions to understand and illustrate empirically the appealing features of LLMs as data generators in the low-data regime, and how our design choices unlock them. We take the Brazilian Covid-19 dataset (Baqui et al., 2020) as a running example and focus on GPT-4 as the LLM.
@@
[Uncaptioned image] Figure 3: Subgroups with fewest samples in $D_{\mathrm{train}}$ benefit the most from data augmentation, on average. Table 1: Deep dive into the top 5 demographic subgroups in the Covid dataset with the largest gains, across 10 seeds, for $|D_{\mathrm{train}}|=20$. GPT-4 improves performance on the smallest groups. Subgroup $n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ Avg. Acc. Gain v. $D_{\mathrm{train}}$ (min - max) GPT-4 TVAE Age_40 0-6 6.38 +- 2.09 -3.37 +- 2.86 Liver 0-1 3.85 +- 3.37 -13.1 +- 3.38 Renal 0-3 4.52 +- 2.01 -18.0 +- 3.22 Amarela 0-1 8.71 +- 1.40 -2.03 +- 2.88 Parda 3-11 5.07 +- 1.50 -6.57 +- 1.61
@@
[Uncaptioned image] Figure 4: Contextual information in the prompt is important for extrapolation. Table 2: Including contextual information in the prompt improves precision (P), recall (R), and utility (U) in low-sample settings (results shown for the Covid dataset). $n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ GPT-4 w/ context GPT-4 no context TVAE P R U P R U P R U 20 $\bf 0.41_{(0.04)}$ $\bf 0.87_{(0.03)}$ $\bf 0.74_{(0.01)}$ $0.13_{(0.0)}$ $0.82_{(0.01)}$ $0.66_{(0.01)}$ $0.33_{(0.07)}$ $0.50_{(0.03)}$ $0.59_{(0.02)}$ 40 $\bf 0.40_{(0.01)}$ $\bf 0.91_{(0.01)}$ $\bf 0.76_{(0.0)}$ $0.11_{(0.0)}$ $0.89_{(0.0)}$ $0.69_{(0.0)}$ $0.27_{(0.01)}$ $0.68_{(0.01)}$ $0.62_{(0.03)}$ 100 $\bf 0.42_{(0.01)}$ $0.86_{(0.02)}$ $\bf 0.75_{(0.01)}$ $0.11_{(0.01)}$ $\bf 0.90_{(0.01)}$ $0.74_{(0.01)}$ $0.39_{(0.02)}$ $0.67_{(0.03)}$ $0.64_{(0.06)}$ 200 $0.44_{(0.02)}$ $0.85_{(0.02)}$ $\bf 0.75_{(0.0)}$ $0.08_{(0.01)}$ $\bf 0.90_{(0.0)}$ $0.60_{(0.01)}$ $\bf 0.47_{(0.0)}$ $0.73_{(0.01)}$ $0.65_{(0.02)}$

### v3

# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in low-data regimes

@@ Abstract
Machine Learning (ML) in low-data settings remains an underappreciated yet crucial problem. Hence, data augmentation methods to increase the sample size of datasets needed for ML are key to unlocking the transformative potential of ML in data-deprived regions and domains. Unfortunately, the limited training set constrains traditional tabular synthetic data generators in their ability to generate a large and diverse augmented dataset needed for ML tasks. To address this challenge, we introduce CLLM, which leverages the prior knowledge of Large Language Models (LLMs) for data augmentation in the low-data regime. However, not all the data generated by LLMs will improve downstream utility, as for any generative model. Consequently, we introduce a principled curation mechanism, leveraging learning dynamics, coupled with confidence and uncertainty metrics, to obtain a high-quality dataset. Empirically, on multiple real-world datasets, we demonstrate the superior performance of CLLM  in the low-data regime compared to conventional generators. Additionally, we provide insights into the LLM generation and curation mechanism, shedding light on the features that enable them to output high-quality augmented datasets.
LLM, Synthetic data, data-centric AI
[…]
@@ 1 Introduction
No data, No Machine Learning. Machine learning (ML) has transformed numerous industries, but its wider adoption is hindered by a pervasive roadblock: insufficient data. Specifically, the use of ML algorithms presumes the availability and access to large datasets for training, be it labeled or unlabeled. Unfortunately, real-world domains are often data scarce: (i) in healthcare and finance, collecting annotations can be expensive or practically impossible; (ii) in developing and low-to-middle income countries (LMICs), digital infrastructure (such as electronic healthcare records (EHRs)) can be limited or nonexistent (Ade-Ibijola & Okonkwo, 2023; Asiedu et al., 2023; Owoyemi et al., 2020; Mollura et al., 2020; Alami et al., 2020; Ciecierski-Holmes et al., 2022) and (iii) within large datasets, there can be (ethnic) minorities that are underrepresented. This lack of data has serious consequences: to sideline these settings to the peripheries of ML advancements and prevent the development of accurate models. How can we build a reliable ML model in this low-data regime, with so few samples? Solving this problem is a major opportunity that would unlock the potential of ML across society, domains, and regions.
Aim. To address this important yet undervalued low-data problem, we aim to augment the small labeled dataset ($n<100$) with synthetic samples. We focus on tabular data, as defining augmentations is non-trivial and can easily result in nonsensical or invalid samples. Moreover, tabular domains like healthcare are often where data scarcity is acute.
@@
Related work. Data augmentation is a widely used and different approach to address data scarcity in tabular data contexts. Methods are either based on generative models (Ghosheh et al., 2023; Biswas et al., 2023; Wang & Pai, 2023; Machado et al., 2022; Tanaka & Aranha, 2019) such as GANs (Xu et al., 2019), VAEs (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), Score-based models (Kotelnikov et al., 2022; Kim et al., 2022), or alternatively traditional methods such as SMOTE (Chawla et al., 2002; Wang & Pai, 2023; Machado et al., 2022). However, in low-data regimes ($n<100$), the training data may not describe the full data distribution well, despite it being i.i.d. draws. Consequently, this harms conventional methods since the augmented data may not be sufficiently diverse and accurate, restricting the generalizability of predictive models trained on such data.
Recent work has shown the potential of fine-tuning Large Language Models (LLMs) for tabular data generation (Borisov et al., 2023). While LLMs offer some degree of prior knowledge, there are two challenges in our setting. First, it is computationally expensive to fine-tune LLMs, while needing specialized hardware —luxuries often not available in LMICs, thereby limiting applicability in such settings. Second, fine-tuning often assumes a large number of samples. In our low-data setting, it could lead to overfitting and low-quality generated samples, and hence poor downstream models—as we show for (Borisov et al., 2023) in Sec. 3. Tangential to data augmentation, prior work has tackled data scarcity in the tabular setting via the lens of transfer learning or few-short learning, by using a knowledge graph (Margeloiu et al., 2022; Ruiz et al., 2023) (which might not be available) or a pretrained model (Levin et al., 2022; Jin & Ucar, 2023; Hegselmann et al., 2023). However, unlike data augmentation, these approaches are not flexible, as they tie the data customer to use a certain downstream predictor. We provide an extended discussion on this point in Appendix A.
@@
Curated LLMs. To address the shortcomings of the aforementioned augmentation approaches, we propose Curated LLM (CLLM). First, CLLM  leverages the in-context capabilities of LLMs for generation, thereby reducing the computational burden compared to fine-tuning. We also posit for the low-data regime; the diverse pretraining corpus of LLMs carries valuable prior knowledge, which may offer more diversity in their generation compared to other conventional tabular generators. Of course, LLMs are not perfect. Consequently, balancing the utility of LLMs against the risk of noisy, irrelevant data is important to ensure reliable downstream performance. Hence, this necessitates systematic assessment of the generated data. In fact, this issue is vital for any generative model.
@@
This motivates the second key aspect of CLLM, i.e. a post-generation data curation mechanism. This addresses the overlooked aspect that not all of the synthetic samples are useful to downstream model performance, with some samples even harmful. We anchor our approach with ideas from learning theory that show the behavior of individual data samples during training, called learning dynamics, provides a salient signal about the value of samples to a learner (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). To provide intuition, samples with variable predictions might be considered ambiguous or other samples might never be learned correctly and could harm a model. In CLLM, we study the learning dynamics of the synthetic data samples, with respect to a model trained on the small real dataset. We then analyze these dynamics by computing two key metrics: confidence and aleatoric (data) uncertainty. These metrics form the basis for curating the synthetic samples. We then aim to enable a highly performant downstream model when trained on the curated dataset.
{mdframed}
@@
[leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, skipbelow=0pt] Contributions: CLLM  is a novel data augmentation approach allying the strengths of LLMs with a robust data curation mechanism to improve data augmentation in the low-data regime $(n<100)$, bringing several contributions: Improved performance: we empirically demonstrate on 7 real-world datasets that CLLM  enables superior downstream performance compared to 6 widely used tabular data generative models and data augmentation techniques. Value of curation: we show the overlooked aspect of synthetic data curation improves downstream performance across the generative models. This highlights the flexibility and broad utility of our curation mechanism for data augmentation. Insights: we dissect the two aspects of CLLM (LLM and data curation) along a variety of dimensions, providing insights and understanding into why the approach is beneficial. We show the largest gains are for underrepresented subgroups and in low-data settings. These contributions, which address the data logjam, pave the way towards wider usage of ML across society, domains and regions.
@@ 2 CLLM: Synergy of LLM Generation and Data Curation
Set-up. Given feature space $\mathcal{X}$, and label space $\mathcal{Y}=\{1,...,k\}$, we assume that we only have a small labeled dataset $D_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, with $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ and $n<100$ (low data setting). Assume $D_{\mathrm{train}}$ is drawn i.i.d. from the real distribution $p_{R}(X,Y)$. We also assume access to a pretrained LLM to generate samples. We denote the output distribution of the LLM as $p_{\Phi}(X,Y)$, with $\Phi$ containing parameters that we control (e.g., input prompts). Our goal is to generate a dataset to augment the small $D_{\mathrm{train}}$, and subsequently use it to train a classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$. Successful augmentation will provide a better classifier $f$, than if we had trained $f$ on the small $D_{\mathrm{train}}$ itself. We measure downstream performance on a separate held-out dataset of real data, $D_{\mathrm{test}}$.
@@
In each sub-section we describe and motivate the design of the different aspects of CLLM  (LLM and curation mechanism). Furthermore, we provide insights and understanding into their role in improving data utility, which we later quantify on multiple real-world datasets in Sec. 3.
@@ 2.1 Data generation with LLMs based on a small $D_{\mathrm{train}}$
As outlined in Sec. 1, in the low-data regime, conventional tabular generative models (e.g. CTGAN, TVAE) are constrained by the limited $D_{\mathrm{train}}$ and may not generate sufficiently diverse and/or accurate synthetic data. To address this, we propose to leverage LLMs, building on their large-scale pretraining. We first outline the appealing properties of LLMs for tabular data generation when we have very few samples, then describe design choices to exploit these.
  • Prior knowledge. LLMs are pretrained with a vast corpus of information (Chowdhery et al., 2022; Singhal et al., 2023). When prompted to generate samples with limited real data, LLMs can leverage this encoded prior information about similar problems and feature-label relationships to enhance both accuracy and diversity of generation.
@@
Motivation for a frozen LLM. Using a frozen black-box LLM (e.g. GPT-4 or GPT-3.5) is computationally cheaper and requires less specialized hardware (i.e. GPUs) compared to fine-tuning. This relates to settings described in Sec. 1, such as LMICs, where we may not have the computational resources to fine-tune an LLM. Even in settings where fine-tuning is possible, we show empirically in Sec. 3 that LLM fine-tuning (e.g. GReaT baseline) is suboptimal in low-data settings ($n<100$) compared to providing in-context examples coupled with curation.
Dissecting the LLM’s generative features. We now investigate various dimensions to understand and illustrate empirically the appealing features of LLMs as data generators in the low-data regime, and how our design choices unlock them. We use the Brazil Covid-19 dataset (Baqui et al., 2020) as a running example and focus on GPT-4 as the LLM.
@@
Figure 3: Subgroups with fewest samples in $D_{\mathrm{train}}$ benefit the most from data augmentation, on average.
Fig. 3 shows the performance gain obtained by training a classifier on data generated by GPT-4 compared to training on the small $D_{\mathrm{train}}$. The greatest gains, on average, are for subgroups for which we have no data in $D_{\mathrm{train}}$, yet GPT-4 can extrapolate and generate samples for these subgroups. This further validates the rationale of extrapolation via prior knowledge as a key source of gain for GPT-4.
Table 1 shows fine-grained results (across 10 different seeds) for the 5 subgroups that benefit the most from data augmentation, which are small-sized demographic subgroups. This finding has real-world implications for equity, showing we can improve performance for underrepresented subgroups even when we lack data or collecting data is difficult/costly.
Table 1: Deep dive into the top 5 demographic subgroups in the Covid dataset with the largest gains, across 10 seeds, for $|D_{\mathrm{train}}|=20$. GPT-4 improves performance on the smallest groups.
Subgroup $n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ Avg. Acc. Gain v. $D_{\mathrm{train}}$
(min - max) GPT-4 TVAE
Age_40 0-6 6.38 +- 2.09 -3.37 +- 2.86
Liver 0-1 3.85 +- 3.37 -13.1 +- 3.38
Renal 0-3 4.52 +- 2.01 -18.0 +- 3.22
Amarela 0-1 8.71 +- 1.40 -2.03 +- 2.88
Parda 3-11 5.07 +- 1.50 -6.57 +- 1.61
@@
Figure 4: Contextual information in the prompt is important for extrapolation.
Fig. 4 qualitatively shows that not including contextual knowledge in the prompt gives lower coverage of $D_{\mathrm{oracle}}$ with less extrapolation beyond $D_{\mathrm{train}}$. We quantify this in Table 2 using Precision (Quality) and Recall (Diversity) metrics (Sajjadi et al., 2018), as well as Utility (Downstream performance). GPT-4 with contextual information has superior precision and recall in the low data setting. Furthermore, we show that the lack of contextual information in the prompt significantly harms the precision (quality) of the data even compared to TVAE. This highlights that LLMs need guidance, as we are only able to get the extrapolation and performance benefits by including contextual information, further motivating our design choices in the prompt. We conduct the same experiment with the Compas dataset in Appendix C.3.
Table 2: Including contextual information in the prompt improves precision (P), recall (R), and utility (U) in low-sample settings (results shown for the Covid dataset).
P R U P R U P R U
20 $\bf 0.41_{\scriptsize(0.04)}$ $\bf 0.87_{\scriptsize(0.03)}$ $\bf 0.74_{\scriptsize(0.01)}$ $0.13_{\scriptsize(0.0)}$ $0.82_{\scriptsize(0.01)}$ $0.66_{\scriptsize(0.01)}$ $0.33_{\scriptsize(0.07)}$ $0.50_{\scriptsize(0.03)}$ $0.59_{\scriptsize(0.02)}$
40 $\bf 0.40_{\scriptsize(0.01)}$ $\bf 0.91_{\scriptsize(0.01)}$ $\bf 0.76_{\scriptsize(0.0)}$ $0.11_{\scriptsize(0.0)}$ $0.89_{\scriptsize(0.0)}$ $0.69_{\scriptsize(0.0)}$ $0.27_{\scriptsize(0.01)}$ $0.68_{\scriptsize(0.01)}$ $0.62_{\scriptsize(0.03)}$
100 $\bf 0.42_{\scriptsize(0.01)}$ $0.86_{\scriptsize(0.02)}$ $\bf 0.75_{\scriptsize(0.01)}$ $0.11_{\scriptsize(0.01)}$ $\bf 0.90_{\scriptsize(0.01)}$ $0.74_{\scriptsize(0.01)}$ $0.39_{\scriptsize(0.02)}$ $0.67_{\scriptsize(0.03)}$ $0.64_{\scriptsize(0.06)}$
200 $0.44_{\scriptsize(0.02)}$ $0.85_{\scriptsize(0.02)}$ $\bf 0.75_{\scriptsize(0.0)}$ $0.08_{\scriptsize(0.01)}$ $\bf 0.90_{\scriptsize(0.0)}$ $0.60_{\scriptsize(0.01)}$ $\bf 0.47_{\scriptsize(0.0)}$ $0.73_{\scriptsize(0.01)}$ $0.65_{\scriptsize(0.02)}$